# Human-Activity AGV Quality Assessment: A Benchmark Dataset and an Objective Evaluation Metric

### v1

@@ 9.2.3 Action Quality Assessment Methods
CoRe [97] evaluates action quality by regressing relative scores against a reference video with similar attributes, such as action category. Due to differences in categorization strategies between the Human-AGVQA and AQA-7 datasets, we randomly selected an exemplar video generated by another T2V model within the same prompt category. Each video was divided into four equal segments, each containing four consecutive frames.
TSA [92] is a temporal segmentation attention module designed for procedure-aware cross-attention learning after spatial-temporal feature extraction. Similar to CoRe, each video is evenly divided into segments of four consecutive frames and processed using the I3D model.
For Motion Smoothness, Temporal Flickering, Action-Score, and Flow-Score metrics, we directly used their respective implementation code in VBench [33] and EvalCrafter [54] without specific changes.
9.3 Training details
@@
9.4 Ablation Study: The Effectiveness of Different Backbones
@@
9.4.1 Body-part Segmentation Network
@@
Table 9: Ablation Study of Body Segmentation Network.
@@
None 0.659 0.671 0.692 0.704 0.745 0.752 0.581 0.524
OpenPose [4] 0.757 0.764 0.740 0.748 0.753 0.751 0.745 0.667
Deeplabv3+ [10] 0.712 0.725 0.724 0.736 0.758 0.767 0.649 0.546
JPPNet [50] 0.685 0.692 0.726 0.732 0.773 0.782 0.582 0.475
HRNet [69] 0.772 0.783 0.704 0.714 0.763 0.773 0.675 0.588
Sapiens [36] 0.812 0.821 0.788 0.792 0.780 0.788 0.815 0.752
Table 10: Ablation Study of Text Feature Extraction Network.
@@
None 0.781 0.789 0.754 0.765 0.773 0.776 0.802 0.736
CLIPScore [27] 0.795 0.803 0.758 0.764 0.778 0.778 0.810 0.737
BLIP [45] 0.806 0.812 0.772 0.782 0.768 0.773 0.810 0.744
viCLIP [85] 0.803 0.807 0.769 0.773 0.767 0.774 0.813 0.739
ImageReward [93] 0.801 0.805 0.775 0.778 0.772 0.779 0.811 0.735
HPSv2 [91] 0.810 0.817 0.779 0.790 0.780 0.787 0.814 0.751
PickScore [37] 0.812 0.821 0.788 0.792 0.780 0.788 0.815 0.752
9.4.2 Holistic Quality Feature Extraction Network
@@
Table 11: Ablation Study of Holistic Quality Feature.
@@
None 0.704 0.722 0.681 0.695 0.727 0.734 0.761 0.710
CLIPScore [27] 0.728 0.739 0.712 0.719 0.743 0.748 0.772 0.717
viCLIP [85] 0.723 0.731 0.723 0.733 0.732 0.744 0.789 0.731
CLIP-B/16 (OpenAI) [64] 0.726 0.735 0.717 0.728 0.738 0.749 0.792 0.735
CLIP-B/32 (OpenAI) [64] 0.739 0.745 0.731 0.742 0.746 0.752 0.787 0.729
CLIP-L/14 (OpenAI) [64] 0.752 0.757 0.743 0.749 0.751 0.766 0.798 0.740
CLIP-L/14 (LAION) [12] 0.776 0.785 0.768 0.777 0.760 0.768 0.802 0.743
CLIP-H/14 (LAION) [12] 0.799 0.812 0.785 0.791 0.772 0.780 0.811 0.745
CLIP-G/14 (LAION) [12] 0.795 0.805 0.782 0.786 0.768 0.778 0.805 0.737
HPSv2 [91] 0.809 0.814 0.788 0.790 0.779 0.784 0.814 0.750
PickScore [37] 0.812 0.821 0.788 0.792 0.780 0.788 0.815 0.752
Table 12: Ablation Study of Action Quality Analyzer.
@@
None 0.767 0.770 0.747 0.748 0.728 0.740 0.753 0.704
I3D [5] 0.781 0.789 0.752 0.762 0.709 0.713 0.811 0.739
C3D [76] 0.797 0.798 0.760 0.773 0.709 0.713 0.811 0.739
X3D [18] 0.808 0.812 0.784 0.790 0.764 0.769 0.813 0.747
SlowFast [19] 0.812 0.821 0.788 0.792 0.780 0.788 0.815 0.752
9.4.3 Text Feature Extraction Network
@@
9.4.4 Action Quality Analyzer

### v2

@@ C.3.3. Action Quality Assessment Methods
CoRe (Yu et al., 2021) evaluates action quality by regressing relative scores against a reference video with similar attributes, such as action category. Due to differences in categorization strategies between the Human-AGVQA and AQA-7 datasets, we randomly selected an exemplar video generated by another T2V model within the same prompt category. Each video was divided into four equal segments, each containing four consecutive frames.
TSA (Xu et al., 2022) is a temporal segmentation attention module designed for procedure-aware cross-attention learning after spatial-temporal feature extraction. Similar to CoRe, each video is evenly divided into segments of four consecutive frames and processed using the I3D model.
For Motion Smoothness, Temporal Flickering, Action-Score, and Flow-Score metrics, we directly used their respective implementation code in VBench (Huang et al., 2024) and EvalCrafter (Liu et al., 2024a) without specific changes.
C.4. Training details
@@
C.5. Ablation Study: The Effectiveness of Different Backbones
@@
C.5.1. Body-part Segmentation Network
@@
Table 9. Ablation Study of Body Segmentation Network.
@@
None 0.651 0.666 0.685 0.701 0.742 0.746 59.31 53.22
OpenPose (Cao et al., 2017) 0.749 0.759 0.738 0.744 0.745 0.747 75.58 67.74
Deeplabv3+ (Chen et al., 2018) 0.706 0.718 0.719 0.731 0.754 0.760 66.33 56.05
JPPNet (Liang et al., 2018) 0.677 0.686 0.722 0.726 0.759 0.768 59.55 48.68
HRNet (Sun et al., 2019) 0.767 0.775 0.699 0.706 0.757 0.764 68.91 59.94
Sapiens (Khirodkar et al., 2024) 0.805 0.809 0.771 0.778 0.768 0.773 83.02 75.94
Table 10. Ablation Study of Text Feature Extraction Network.
@@
None 0.774 0.779 0.737 0.742 0.759 0.760 80.53 74.16
CLIPScore (Hessel et al., 2021) 0.791 0.795 0.741 0.748 0.763 0.765 81.28 74.11
BLIP (Li et al., 2022a) 0.798 0.805 0.760 0.765 0.754 0.755 81.43 74.75
viCLIP (Wang et al., 2022) 0.799 0.799 0.754 0.761 0.751 0.758 81.71 74.23
ImageReward (Xu et al., 2023) 0.797 0.798 0.763 0.768 0.759 0.764 81.66 74.09
HPSv2 (Wu et al., 2023a) 0.802 0.808 0.762 0.771 0.765 0.771 81.62 75.23
PickScore (Kirstain et al., 2023) 0.805 0.809 0.771 0.778 0.768 0.773 83.02 75.94
C.5.2. Holistic Quality Feature Extraction Network
@@
Table 11. Ablation Study of Holistic Quality Feature.
@@
None 0.696 0.713 0.666 0.682 0.709 0.717 76.61 71.49
CLIPScore (Hessel et al., 2021) 0.719 0.729 0.699 0.702 0.730 0.734 77.78 72.21
viCLIP (Wang et al., 2022) 0.713 0.723 0.710 0.719 0.720 0.727 79.43 73.63
CLIP-B/16 (OpenAI) (Radford et al., 2021b) 0.718 0.726 0.703 0.712 0.723 0.732 79.87 73.97
CLIP-B/32 (OpenAI) (Radford et al., 2021b) 0.732 0.739 0.714 0.726 0.731 0.735 79.22 73.54
CLIP-L/14 (OpenAI) (Radford et al., 2021b) 0.743 0.748 0.729 0.736 0.736 0.750 80.27 74.22
CLIP-L/14 (LAION) (Cherti et al., 2023) 0.766 0.779 0.751 0.761 0.747 0.752 80.54 74.55
CLIP-H/14 (LAION) (Cherti et al., 2023) 0.793 0.804 0.769 0.774 0.758 0.764 81.18 75.01
CLIP-G/14 (LAION) (Cherti et al., 2023) 0.787 0.795 0.767 0.769 0.755 0.761 80.81 74.13
HPSv2 (Wu et al., 2023a) 0.801 0.805 0.771 0.776 0.765 0.767 81.39 75.24
PickScore (Kirstain et al., 2023) 0.805 0.809 0.771 0.778 0.768 0.773 83.02 75.94
Table 12. Ablation Study of Action Quality Analyzer.
@@
None 0.751 0.762 0.734 0.735 0.712 0.723 76.15 70.72
I3D (Carreira and Zisserman, 2017b) 0.764 0.777 0.735 0.748 0.695 0.701 81.83 73.73
C3D (Tran et al., 2015) 0.782 0.788 0.748 0.755 0.696 0.699 81.71 74.18
X3D (Feichtenhofer, 2020) 0.793 0.801 0.767 0.774 0.750 0.757 81.96 74.84
SlowFast (Feichtenhofer et al., 2019) 0.805 0.809 0.771 0.778 0.768 0.773 83.02 75.94
C.5.3. Text Feature Extraction Network
@@
C.5.4. Action Quality Analyzer